40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Thu Sep 4 19:39:36 2014, Tara, Evan, DailyProgress, FSS, TTFSS OLTFs pdhTFs.pdf
    Reply  Sun Sep 7 11:54:21 2014, Evan, DailyProgress, FSS, TTFSS OLTFs, residual frequency noise pdhTFs.pdfoltfNorth.pdfoltfSouth.pdffrnoise.pdfnoiseBudget.pdf
       Reply  Mon Sep 8 04:20:41 2014, rana, DailyProgress, FSS, TTFSS OLTFs, residual frequency noise 
       Reply  Wed Oct 2 19:09:31 2019, anchal, DailyProgress, FSS, TTFSS OLTFs; Repeated this 5 year old measurement FSS_OLTF.pdf
          Reply  Thu Jan 9 19:00:19 2020, anchal, DailyProgress, FSS, TTFSS OLTFs FSS_OLTF.pdf
             Reply  Wed Jan 22 14:16:28 2020, anchal, DailyProgress, FSS, TTFSS OLTFs FSS_OLTF.pdf
                Reply  Wed Jan 22 16:00:26 2020, anchal, DailyProgress, FSS, TTFSS OLTFs with Maximum Gain FSS_OLTF.pdf
                   Reply  Thu Jan 23 19:23:03 2020, anchal, DailyProgress, FSS, NFSS Boost Stage NorthFSSBosstStageTF.pdf
Message ID: 2443     Entry time: Wed Oct 2 19:09:31 2019     In reply to: 1504     Reply to this: 2508
Author: anchal 
Type: DailyProgress 
Category: FSS 
Subject: TTFSS OLTFs; Repeated this 5 year old measurement 

I repeated the measurement done by Tara and Evan to update residual NPRO noise in our noise budget.

Quote:

Acutally it does look like it's a 50 Ω loading issue. I find that when 50 Ω inline terminators are added to OUT1 and OUT2, the measured OLTF is reduced by a factor of 1.6. This explains the discrepancy between the SR785 and HP4395 measurements. I've attached the corrected OLTF plots, along with plots of a vector fit, and the expected residual frequency noise [assuming a free-running NPRO noise of 104 Hz/Hz1/2 × (1 Hz / f)].

South UGF is at 200 kHz with almost no phase margin. We need to fix this.


Conclusions:

  • Clearly, the North side's boost isn't working as was reported in CTN:2384
  • South side looks as healthy as North in CTN:1504
  • Tomorrow I'll update the noisebudget with new residual NPRO noise, for now, the second plot gives the prediction.
  • It matches very well with our observed floor of noise in 100Hz to 1 kHz region in the beatnote spectrum.

Code and Data

Attachment 1: FSS_OLTF.pdf  91 kB  Uploaded Wed Oct 2 20:39:23 2019  | Hide | Hide all
FSS_OLTF.pdf FSS_OLTF.pdf
ELOG V3.1.3-